# Is JPEG AI going to change image forensics?

### v1

@@ 5.1 Deepfake image detection
Tab. 4 depicts the achieved detection metrics over pristine vs. synthetic (first column) and pristine vs. JPEG AI samples (other columns) before and after retraining and fine-tuning. The detector is still capable of classifying real vs. synthetic samples, with a small performance loss for the retraining case and, as could be expected, a larger drop when fine-tuning. Indeed, the fine-tuning setup did not include uncompressed versions of pristine and synthetic data; thus, we could expect to lose generalization slightly.
@@
We can notice a reduction of the JPEG AI counter-forensic effect as all metrics are reduced. However, the AUC is still above for BPPs lower than , and the false alarms (i.e., TPR) progressively reduce but are still for BPPs lower than . These preliminary results show that a huge augmentation campaign should probably be considered to be completely robust to this compression.
@@ 5.2 Image splicing localization
Visual inspection. As introduced in Sec. 4, we analyze two detectors, TruFor and MMFusion. We start by visually inspecting the output of both on a sample from the DSO1 dataset. Figure 7 shows the output masks for the original format as well as the masks obtained after compressing the input sample with JPEG AI at different BPPs. We display the original spliced sample together with its compressed versions as well.
@@
Even though the quality of the JPEG AI samples is visually high, we can observe that the compression operation heavily degrades the output masks in two different ways. First, for low BPP, the original manipulated area, while still largely identified, may present several false negatives, i.e., manipulated pixels that are incorrectly classified as pristine. Second, we can observe many false positives in the rest of the image, i.e., pristine pixels erroneously classified as manipulated at all BPP values. These two separate phenomena suggest that JPEG AI can contemporarily “hide” some manipulation traces while introducing artifacts mistaken for tampered with pixels.
| Metric | Original | JPEG AI compressed (BPP) | ||||||
| [38] | 0.57 | 0.64 | 0.73 | |||||
| @0.5 | 0.55 | 0.59 | 0.67 | |||||
| best | 0.58 | 0.63 | 0.71 | 0.71 | ||||
| F1@0.5 | 0.36 | 0.42 | 0.53 | |||||
| F1 best | 0.49 | 0.56 | 0.66 | |||||
| FPR @0.5 | 0.15 | 0.18 | 0.23 | 0.23 | 0.23 | 0.15 | ||
| [78] | 0.71 | 0.75 | 0.82 | |||||
| @0.5 | 0.60 | 0.67 | 0.74 | |||||
| best | 0.67 | 0.71 | 0.77 | |||||
| F1@0.5 | 0.37 | 0.43 | 0.57 | |||||
| F1 best | 0.48 | 0.56 | 0.68 | |||||
| FPR @0.5 | 0.11 | 0.10 | 0.09 | 0.09 | ||||
Quantitative analysis. We compute the pixel-level localization metrics illustrated in Sec. 4 to quantify this visual intuition and report their values in Tab. 5. In particular, we calculate these metrics for all the datasets considered and then average their values2. As the reader can notice, performance drops sharply as the compression ratio increases. For TruFor, we have a drop of points in AUC for the compression at BPP . This result indicates that it is harder to find a correct threshold to impose on to classify most pixels correctly. We motivate this behavior by a higher number of false alarms, as highlighted from the visual inspection of Figure 7, and from the FPR metric that shows an increase of at least points for all BPP values.
We also see a sharp decrease in BA and F1-score, evaluated at two standard thresholds, the one returning the best values for each image and the common threshold . It is interesting to notice that, in high BPP settings, both F1 and BA scores still show a drop of . Therefore, even if not visible to the naked eye, we conjecture that interpolation artifacts are still present and can impair the localization performance of the detectors considered.
We can notice a comparable trend in MMFusion’s results as well. There is a decrease in AUC in the order of to points for all BPPs; an increase of FPR from to in the BPP scenario and a decrease between to and from to for BA and F1-score, respectively, for both thresholds considered. Therefore, we can formulate similar considerations to TruFor, even though better robustness is noticeable, which, in this case, might be connected to the MMFusion combination of different detectors’ outputs.
Finally, we refer the reader to Supplementary Material Sec. 9 for the effects of standard JPEG. JPEG has less impact on the detectors’ performance for all quality factors considered. We argue that while JPEG is known for “canceling” or “hiding” tampering traces, JPEG AI, as discussed, has a heavier effect due to the various upsampling artifacts introduced.
@@ 6 Discussion and conclusion
While far from being an exhaustive and complete analysis of the forensic implications of adopting JPEG AI, our results highlight how off-the-shelf forensic detectors need to be prepared to handle the artifacts of this upcoming standard. In both deepfake image detection and image splicing localization tasks, heavy compression leads to inaccurate classification of pristine content as manipulated, requiring practitioners to consider JPEG AI images in their work.
@@ 7 Experimental setup additional details
Experiments reproducibility. All code for replicating our experiments is available at the following URL https://anonymous.4open.science/r/jpeg-ai-antifor-4818/README.md.
@@
[59]: an EfficientNet-B4 [76] performing a patch-based analysis of the image and aggregating multiple scores to decide the “fakeness” of the sample. This detector is trained exclusively on datasets of human faces, with deepfake samples coming from various generators, including GANs, transformers, and diffusion models.
Deepfake datasets. Regarding the preliminary analysis of Sec. 5.1, to test the performances of every detector in normal conditions we include for every pristine test set a deepfake “counterpart”. These images have been selected by the original authors of the detectors to depict similar content to the pristine one, but are generated using different deepfake image generation techniques. In total, we considered different generation techniques ranging from GANs to diffusion models, namely Guided Diffusion [33], Latent Diffusion [70], Taming Transformers [36], DALL·E [69] in various versions (i.e., mini [2], 2 [7], and 3 [8]), GLIDE [64], Stable Diffusion [70] in various versions (i.e., 1.3 [12], 1.4 [13], XL [15], and 2 [14]), Adobe Firefly [9], MidJourney v5 [11], BigGAN [23], GauGAN [67], StyleGAN [49], StyleGAN2 [51], ProGAN [48], and StarGAN [26]. Tab. 6 report their relationships with the different pristine datasets used in our evaluation.
@@ 8 Deepfake image detection additional results
Preliminary analysis on synthetic detection. Tab. 7 reports the results for the simple case of synthetic image detection without JPEG AI image compression of the pristine samples. We report the complete numbers for all detectors on all datasets.
Please notice that we tested each detector only on the datasets on which the original authors first proposed the method. The only exception is represented by [59] since the authors reported good generalization capabilities on various semantic content despite training their solution only on human faces. As the reader can quickly inspect, all detectors achieve acceptable performance, with only two cases below the threshold for the AUC, i.e., [29]-B on LAION and [59] on COCO.
@@
Scores distributions. Fig. 8 to Fig. 11 report the scores distribution of the detectors [37], [29]-A, [65], and [59] while varying the JPEG AI compression ratio or the JPEG quality factor of the compressed images. We report one separate plot for each of the datasets on which we tested the detectors. Results are in line with the ones reported in Fig. 4 and Fig. 5 of the main paper: while increasing the JPEG AI compression ratio creates a “shift” of the scores distribution towards the right, i.e., towards “deepfake”-like scores, standard JPEG does not present the same behavior regardless of the strength used for saving the data. We refer the reader to the code repository to visualize more distribution score plots.
Complete analysis of JPEG AI compression ratio. Tab. 9 to Tab. 15 report the results of the experiments by varying the compression ratio of JPEG AI for each detector on every single dataset. For all tables, as commonly done in deepfake detection tasks, we evaluate the BA and TPR at threshold.
| Synthetic | BPP = 0.12 | BPP = 0.25 | BPP = 0.5 | BPP = 0.75 | BPP = 1.0 | BPP = 2.0 | ||||||||||||||||||||||
| EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | ||||||||||||||||||||||
| LSUN | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
@@
Comparison with standard JPEG. Tab. 16 to Tab. 23 report the results of the experiments varying the quality factor of JPEG for each detector on every single dataset. For all tables, as commonly done in deepfake detection tasks, we evaluate the TPR and BA at the threshold.
| Synthetic | QF = 65 | BPP = 75 | QF = 85 | QF = 95 | ||||||||||||||||
| EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | ||||||||||||||||
| LSUN | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
@@
| LAION | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| FFHQ | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| Synthetic | QF = 65 | BPP = 75 | QF = 85 | QF = 95 | ||||||||||||||||
| EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | ||||||||||||||||
| LSUN | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
@@
| FFHQ | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| Synthetic | QF = 65 | BPP = 75 | QF = 85 | QF = 95 | ||||||||||||||||
| EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | EMD / AUC / BA / TPR | ||||||||||||||||
| LSUN | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| Imagenet | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| RAISE | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| LAION | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||

### v2

@@ 5.1 Deepfake image detection
Tab. 4 depicts the achieved detection metrics over non-JPEG AI data (first left column) and over JPEG AI images (other columns) before and after retraining and fine-tuning. The detector is still capable of classifying original format samples, experiencing a slight performance loss in the retraining case and, as expected, a more significant drop when fine-tuning. Indeed, the fine-tuning setup includes only JPEG AI-compressed samples; thus, a subtle loss of generalization is expected.
@@
As for what regards the BA and the FPR, we can notice a reduction of the JPEG AI counter-forensic effect since these metrics improve upon the initial results. Nonetheless, the BA is still below , and the FPR is above for lower BPPs. Instead, the effects of retraining on the FNR are less visible: as observed before, though the FNR tends to reduce for low BPPs, this number gradually increases, achieving even higher values than those resulting when processing the original format images. These preliminary results show that a huge augmentation campaign should probably be considered to be completely robust to JPEG AI compression.
@@ 5.2 Image splicing localization
Visual inspection. As introduced in Sec. 4, we analyze three detectors, TruFor [43], MMFusion [85], and IFOSN [92]. We start by visually inspecting their output on a sample from the DSO-1 dataset. Figure 6 shows the output masks for the original format as well as the masks obtained after compressing the input sample with JPEG AI at different BPPs for TruFor and MMFusion. We display the original spliced sample together with its compressed versions as well.
@@
Even though the quality of the JPEG AI images is visually high, we can observe that the compression operation heavily degrades the output masks in two different ways. First, for low BPP, the original manipulated area, while still largely identified, may present several false negatives, i.e., manipulated pixels that are incorrectly classified as pristine. Second, we can observe many false positives in the rest of the image, i.e., pristine pixels erroneously classified as manipulated at all BPP values. These two separate phenomena suggest that JPEG AI can contemporarily “hide” some manipulation traces while introducing artifacts mistaken for tampered with pixels.
Quantitative analysis. We compute the pixel-level localization metrics illustrated in Sec. 4 to quantify this visual inspection and report their values in Tab. 5 (second and third right column). In particular, we calculate these metrics for all the datasets considered and then average their values.
As the reader can notice, performance drops sharply as the JPEG AI compression ratio increases for all detectors. For the compression at BPP , we have a drop of more than and points in AUC and BA, respectively. Moreover, the F1 score sharply decreases by more than points.
It is interesting to notice that, in high BPP settings, the F1 score may still show a drop of . Even if not visible to the naked eye, we conjecture that JPEG AI interpolation artifacts can impair the localization performance of the detectors considered. The IFOSN detector [92] is the least affected detector in case of high BPPs: we believe this might be due to its specific training process, developed for gaining robustness against social network compression.
Comparison with standard JPEG. The last row of Figure 6 shows the output masks obtained after compressing the input sample with JPEG at (i.e., the worst considered QF in our experiments) for TruFor [43] and MMFusion [85]. Notice the differences to JPEG AI: in the JPEG case, the localization performances are only slightly reduced, and the tampered with area is still well detectable.
We report in Tab. 5 (last column) the quantitative results for all detectors when tested against JPEG compression. It is worth noticing that JPEG impacts the detectors’ performance much less for all the QFs considered. In particular, even at the lowest QF, all metrics drop by a maximum of points. We argue that, while JPEG is known for “hiding” tampering traces, JPEG AI has a considerably heavier effect due to the various upsampling artifacts introduced.
5.3 Double JPEG AI experiments.
A very common forensic scenario is that represented by double compression [33, 20, 96]. In this situation, the input samples might have undergone several compression stages before or after the manipulation. For instance, the manipulated area in a spliced sample might present traces of a compression operation performed before the splicing. Such traces can be extremely useful for understanding the overall image history and better localizing the tampered with pixels. Moreover, the double compression setup is also valuable for testing image detectors’ robustness, i.e., how good they are when analyzing images undergoing several compression stages.
In our work, we perform a preliminary analysis of the double compression scenario in the case of JPEG AI. To simplify the experimental setting, we consider two opposite setups: i) the original-format images have been JPEG AI-compressed with various BPP values and then compressed with the lowest BPP considered, i.e., BPP ; ii) the original-format images have been JPEG AI-compressed with various BPP values, and then compressed with the highest BPP considered, i.e., BPP . For the sake of space, we refer the reader to the supplementary material for an in-depth discussion. Our findings show that a strong second JPEG AI compression might cancel traces left by the previous compression step, similar to the behavior observed on standard double JPEG compression [20].
@@ 6 Discussion and conclusion
While far from being an exhaustive and complete analysis of the forensic implications of adopting JPEG AI, our results highlight how off-the-shelf forensic detectors need to be prepared to handle the artifacts of this upcoming standard. In both deepfake image detection and image splicing localization tasks, heavy compression leads to inaccurate classification of the investigated content, requiring practitioners to consider JPEG AI images in their work.
@@
7 Neural image compression and JPEG AI specification
Neural image compression basics. In Neural Image Compression (NIC), an input image is transformed into a compact file or bitstream for storage or transmission, which is later used to reconstruct the original image. The basic approach involves mapping an image vector to a latent representation using a convolutional autoencoder. This latent representation is then quantized to and compressed into a bitstream using an entropy coding method such as arithmetic coding. During reconstruction, the received bits are decoded via reverse entropy coding, and the resulting is fed into a CNN decoder to produce the reconstructed image . These models can be trained end-to-end with the help of differentiable proxies for entropy coding.
As pointed out in Sec. 2, many techniques exist to perform such operations. In 2017, Theis et al. [83] and Ballé et al. [18] introduced the first deep-learning approaches for image compression. [83] proposed compressive autoencoders that learn end-to-end optimized lossy image compression, aiming to outperform traditional codecs, especially at lower bitrates. [18] presented the first pipeline that jointly optimizes all components of the compression pipeline, achieving superior rate-distortion performance compared to traditional codecs. Building upon their previous work, in [19], Ballé et al. incorporate a scale hyperprior to model the entropy of the latent representations more accurately, leading to improved compression efficiency. Minnen et al. [66] combine autoregressive models with hierarchical priors to better capture spatial dependencies in images, enhancing the performance of learned image compression systems. Choi et al. [30] present a conditional autoencoder that enables variable-rate image compression within a single model, adjusting compression rates on the fly without retraining. In 2021, Yang et al. [93] proposed slimmable autoencoders that adjust their network width to accommodate different compression rates, making neural compression more adaptable and efficient. Zhu et al. [98] in 2022 presented a unified entropy model using multivariate Gaussian mixtures to better model the distribution of latent representations, improving compression efficiency.
JPEG AI scheme specification. The astonishing results from the literature pushed in 2019 the JPEG Standardization Committee to launch the first image coding specifications based on end-to-end learning [49], namely the JPEG AI initiative. At its core, JPEG AI follows a similar structure to standard neural image compression algorithms. In particular, we refer the reader to Fig. 2 [50] for the overall JPEG AI standard scheme specification. The encoder consists of a NN-based analysis transform that computes the latent information of the input image . Next, is processed by a latent domain prediction module, generating a latent residual , and by a hyperprior encoder, producing a hyperprior , as proposed in the foundational work by Ballé et al. [19]. For entropy decoding, a latent prediction network retrieves , which is finally input into the synthesis transform at the decoder for reconstructing the image .
As explained in Sec. 2, JPEG AI enforces strict compliance requirements on manufacturers as a standardized format, limiting the customization options within the compression pipeline. Although certain architectural modifications are permitted (indicated by the red modules of Fig. 2), such as adjustments to the analysis transform, to the hyperprior encoder network, or retraining for specific content types to improve performance, these changes are restricted. The modified latent space must still function within the official JPEG AI framework to ensure compatibility [50].
@@ 8 Experimental setup additional details
Experiments reproducibility. All code and additional figures and tables for all detectors is available at the following URL https://github.com/polimi-ispl/jpeg-ai-antifor.
@@
[63]: an EfficientNet-B4 [82] performing a patch-based analysis of the image and aggregating multiple scores to decide the “fakeness” of the sample. This detector is trained exclusively on datasets of human faces, with deepfake samples coming from various generators, including GANs, transformers, and diffusion models;
[81]: a method focusing on detecting deepfakes by analyzing structural artifacts introduced during the up-sampling operations in generative networks like GANs and diffusion models, namely Neighboring Pixel Relationships (NPR). These artifacts create local interdependencies among neighboring pixels in synthetic images. NPR captures and characterizes these relationships to identify forgeries effectively.
Deepfake datasets. To test the performances of every detector in normal conditions, we include for every pristine test set a deepfake “counterpart”. These images have been selected by the original authors of the detectors to depict content similar to that of the pristine ones but are generated using different deepfake image generation techniques. In total, we considered different generation techniques ranging from GANs to diffusion models, namely Guided Diffusion [38], Latent Diffusion [76], Taming Transformers [41], DALL·E [75] in various versions (i.e., mini [2], 2 [6], and 3 [7]), GLIDE [69], Stable Diffusion [76] in various versions (i.e., 1.3 [12], 1.4 [13], XL [15], and 2 [14]), Adobe Firefly [8], MidJourney v5 [11], BigGAN [26], GauGAN [72], StyleGAN [53], StyleGAN2 [55], ProGAN [52], and StarGAN [29]. Tab. 6 report their relationships with the different pristine datasets used in our evaluation.
@@ 9 Deepfake image detection additional results
Preliminary analysis on synthetic detection. Tab. 7 reports the results for the simple case of synthetic image detection without JPEG AI image compression of the test samples. We report the complete numbers for all detectors on all datasets.
We tested each detector only on the datasets on which the original authors first proposed the method. [63] represents the only exception since the authors reported good generalization capabilities on various semantic content despite training their solution only on human faces. As the reader can quickly inspect, all detectors achieve acceptable performance, with only two cases below the threshold for the AUC, i.e., [34]-B on LAION and [63] on COCO.
@@
Scores distributions. Fig. 7 to Fig. 13 report the scores distribution of the detectors [42], [34]-A, [70], and [81] while varying the JPEG AI compression ratio or the JPEG quality factor of the compressed images. We report one plot for each dataset on which we tested the detectors. Results are in line with the ones reported in Fig. 4 and Fig. 5 of the main paper: while increasing the JPEG AI compression ratio creates a “shift” of the scores distribution towards the right, i.e., towards “deepfake”-like scores above the threshold, standard JPEG moves the distribution towards the left, i.e., towards “pristine ”-like scores below the threshold. Such results are particularly relevant for the [81] detector. This tool looks for NPR, i.e., the local interdependence between image pixels caused by up-sampling operators. Such a trace seems particularly sensitive to JPEG compression. Indeed, we can see in Fig. 12(d) and Fig. 12(f) how even the score distribution for pristine samples is moved below the distribution of pristine uncompressed images. However, this result should not surprise us, as it is well known in the forensic community that JPEG compression could conceal traces relative to the image history [73]. We refer the reader to the code repository to visualize more distribution score plots.
Complete analysis of JPEG AI compression ratio. Tab. 9 to Tab. 17 report the results of the experiments by varying the compression ratio of JPEG AI for each detector on every single dataset. For all tables, as commonly done in deepfake detection tasks, we evaluate the BA, FPR, and FNR at the threshold.
| Original format | BPP = 0.12 | BPP = 0.25 | BPP = 0.5 | BPP = 0.75 | BPP = 1.0 | BPP = 2.0 | ||||||||||||||||||||||
| AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | ||||||||||||||||||||||
@@
| Imagenet | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
| LAION | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
| LSUN | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
| RAISE | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
| Original format | BPP = 0.12 | BPP = 0.25 | BPP = 0.5 | BPP = 0.75 | BPP = 1.0 | BPP = 2.0 | ||||||||||||||||||||||
| AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | ||||||||||||||||||||||
| Imagenet | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
| COCO | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
| CelebA | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
| LSUN | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
| LAION | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||||
Comparison with standard JPEG. Tab. 18 to Tab. 26 report the results of the experiments varying the quality factor of JPEG for each detector on every single dataset. For all tables, as commonly done in deepfake detection tasks, we evaluate the BA, FPR, and FNR at the threshold.
| Original format | QF = 65 | BPP = 75 | QF = 85 | QF = 95 | ||||||||||||||||
| AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | ||||||||||||||||
@@
| Original format | QF = 65 | BPP = 75 | QF = 85 | QF = 95 | ||||||||||||||||
| AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | ||||||||||||||||
@@
| Original format | QF = 65 | BPP = 75 | QF = 85 | QF = 95 | ||||||||||||||||
| AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | ||||||||||||||||
@@
| Imagenet | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| LAION | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| LSUN | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| RAISE | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| Original format | QF = 65 | BPP = 75 | QF = 85 | QF = 95 | ||||||||||||||||
| AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | AUC / BA / FPR / FNR | ||||||||||||||||
| Imagenet | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| COCO | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| CelebA | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| LSUN | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||
| LAION | / | / | / | / | / | / | / | / | / | / | / | / | / | / | / | |||||